OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and has a slope of 5? y = 5x - 3 y = 5x - 1 y = 5x + 3 y = 5x - 19

OpenStudy (anonymous):

well what is it? plug in points ;)

OpenStudy (anonymous):

y = 5x - 3

OpenStudy (aravindg):

try again ??

OpenStudy (anonymous):

clue what is the point (0,c) on this line ?

OpenStudy (aravindg):

just plugin the point as @RONNCC said see which equation gets satisfied

OpenStudy (anonymous):

y = 5x + 3

OpenStudy (aravindg):

1=5*4+3 1=20+3 1=23 nice :)

OpenStudy (aravindg):

dont tell anyone !! ;)

OpenStudy (anonymous):

was that it ?

OpenStudy (aravindg):

do you think 1=23

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

geometry

OpenStudy (anonymous):

the only way i know is \[y-(y _{1})=m(x-{x _{1})}\] point slope form

OpenStudy (anonymous):

this is algebra

OpenStudy (anonymous):

not geometry

OpenStudy (anonymous):

@ivandelgado just plug in the points

OpenStudy (zehanz):

Why not use @jalan's formula? It gives you an equation of a line through a given point (x1, y1) with a given slope (m). Fits your problem perfectly! So put in 5 for m, and x1=4, y1=1 and see what happens. Of course, you'll have to shuffle this equation a little to see which of the four options it is...

OpenStudy (anonymous):

\[y-(1)={5(x-(4)}\]

OpenStudy (zehanz):

y - 1 = 5(x - 4), yes. Now expand the brackets...

OpenStudy (zehanz):

@ivandelgado: we've done all we could to help you out here, I think it's your turn now...

OpenStudy (anonymous):

y = 5x - 19

OpenStudy (zehanz):

Well done!

OpenStudy (zehanz):

And remember that neat little formula y-y1=m(x-x1)...

OpenStudy (anonymous):

ok